# Augmented Reality for RObots (ARRO): Pointing Visuomotor Policies Towards Visual Robustness

### v2

@@ IV-B ARRO in Simulation
In addition to the real-world experiments, we also evaluate ARRO in simulation on the tasks pick-v2 and sim-pick-v2. Unlike in pick-v1, the cube is red, and we collected the demonstrations automatically both in the real world and in simulation. We investigate whether ARRO can help mitigate the sim-to-real gap [19] and facilitate cross-embodiment policy transfer. Fig. 8 illustrates our experimental setup: the physical setup in our lab using the FR3 robot, its replication in a MuJoCo [50] simulation environment, and a cross-embodiment variant where the FR3 is replaced by a UR5e robot in the same environment.
@@
Figure 7: Performance comparison across two generalist models, Octo and OpenVLA, fine-tuned on the pick-v1 task (see Fig. 3) and evaluated under altered visual conditions.
@@
Figure 8: Experimental setup for the pick task in two environments and one cross-embodiment scenario. Left to right: The real-world FR3 setup, the model of the setup in MuJoCo used for collecting sim-pick-v2, and the UR5e setup for cross-embodiment experiments in MuJoCo.
TABLE I: Retained success rate of Real-to-sim and cross-embodiment experiments of vanilla, masked and ARRO versions of different policies. 100% means the success rate in the new domain is the same as in the original domain. Real-to-sim was trained on the pick-v2 dataset and cross-embodiment on the sim-pick-v2 dataset. See sections IV-B1 and IV-B2 for details.
@@
Policy Vanilla Masked ARRO Vanilla Masked ARRO
DP 0% 0% 0% 0% 80% 99%
Octo 0% 3% 5% 71% 87% 87%
OpenVLA 0% 53% 55% 29% 72% 80%
@@
Figure 9: Normalized mean evaluation reward per step averaged over 100 episodes on the pick cube task in simulation on the FR3 (same embodiment). All models were trained on the pick-v1 real-world dataset. Shaded areas indicate standard deviations.
IV-B1 Real-to-Sim
Since most datasets for robot imitation learning are collected on real-world setups, we examine a real-to-sim rather than the more typical sim-to-real paradigm to assess ARRO’s performance in domain transfer. We trained Diffusion Policy, Octo, and OpenVLA on the pick-v2 dataset and evaluated the task success rates for two settings: The unmodified real-world scene and the replicated scene in the simulation. We evaluated the trained policies in-distribution on the same unchanged real-world setup on 10 episodes (real-to-real success rate) and out-of-distribution on the replicated MuJoCo setup on 100 episodes (real-to-sim success rate). All models achieve decent real-to-real success rates: 0.9, 0.5 and 0.4 for Diffusion Policy, Octo and OpenVLA, respectively. But as shown in Table I, their performance drops to 0% of their real-world performance when evaluated in the real-to-sim setting. With ARRO, however, OpenVLA retains 55% of its real-world performance. Octo does not benefit as much and only achieves 5% of its original success rate. Diffusion Policy does not exhibit any immediate transfer in terms of success rate, which can be attributed, at least in part, to the missing pre-training stage and differences in the vision backbone architecture. In addition to the success rates in the real-to-sim experiments, we computed reward values using a simple distance metric inspired by ManiSkill [51]. The results are shown in Fig. 9. For all policies, both ARRO and the simple black background yield higher rewards than the vanilla baselines.
IV-B2 Cross-Embodiment
@@
To test the cross-embodiment capabilities of ARRO, we train our policies on a simulation dataset, referred to as sim-pick-v2, using the FR3 and test it on a UR5e embodiment. We evaluate both in-distribution on the training environment for 100 episodes (sim-to-sim success rate) and on the new UR5e embodiment for 100 episodes (cross-embodiment success rate). The results are presented in Table I. The diffusion policy’s performance drops to 0 when evaluated on a novel embodiment, which is expected given that it was trained exclusively on the FR3 robot. In contrast, both the masking with a black background and ARRO exhibit only a minor reduction in success rates, retaining most of their performance. This is likely because the embodiment change has minimal impact on the diffusion policy: the visual input remains largely unaffected due to the masking of the robot body, and the control commands are issued in absolute Cartesian space, which remains invariant to the embodiment.
@@
Octo and OpenVLA also exhibit reduced drops in success rates when combined with ARRO and masking. Unlike the Diffusion Policy, their performance does not drop to zero in the vanilla setting for the new embodiment. This can be attributed to their larger and more diverse pre-training datasets, which are explicitly designed to support cross-embodiment generalization. As a result, both models encountered the UR5e robot during training, facilitating better transfer even in simulation settings.
V Conclusion
@@
In this paper, we presented ARRO, a calibration-free visual augmentation pipeline that improves the robustness of visuomotor policies. By segmenting and retaining only task-relevant regions and overlaying them on a structured virtual background, ARRO enhances generalization without retraining or camera calibration. Our experiments in both real-world and simulated settings demonstrate consistent performance gains for task-specific as well as generalist policies. ARRO thus offers a lightweight and scalable solution for robust robotic manipulation in visually diverse environments.
Looking ahead, an interesting line of research is to apply ARRO at scale using data from the Open X-Embodiment dataset collection [13] to explore its potential for enhancing cross-task and cross-embodiment generalization. Another promising direction is the integration of ARRO with reinforcement learning pipelines to assess its benefits in online policy adaptation. Taken together, these future directions may further expand the applicability of ARRO toward more general, robust, and scalable robotic learning systems.
@@ ACKNOWLEDGMENT
The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the BayernKI project no. v108be. BayernKI funding is provided by Bavarian state authorities.

### v3

@@ IV-C ARRO in Simulation
In addition to the real-world experiments, we also evaluate ARRO in simulation on the tasks pick-v2 and sim-pick-v1. Unlike in pick-v1, the cuboid is red, and we collected the demonstrations automatically via a script in both the real world and the simulation. For $\pi_{0}$ we choose pick-wrist-v1, which includes the wrist camera. We investigate whether ARRO can help mitigate the real-to-sim gap [Li.2024] and facilitate cross-embodiment policy transfer. Fig. 8 illustrates our experimental setup: the physical setup in our lab using the FR3 robot, its replication in a MuJoCo [mujoco] simulation environment, and a cross-embodiment variant where the FR3 is replaced by a UR5e robot in the same environment.
IV-C1 Real-to-Sim
Since most datasets for robot imitation learning are collected on real-world setups, we examine a real-to-sim rather than the more typical sim-to-real paradigm to assess ARRO’s performance in domain transfer. We trained Diffusion Policy, Octo and OpenVLA on the pick-v2 dataset, and $\pi_{0}$ on the pick-wrist-v1 dataset, and evaluated the task success rates for two settings: The unmodified real-world scene and the replicated scene in the simulation. We evaluated the trained policies in-distribution on the same unchanged real-world setup on 10 episodes (real-to-real success rate) and out-of-distribution on the replicated MuJoCo setup on 100 episodes (real-to-sim success rate). All models achieve decent real-to-real success probabilities in the vanilla case: 90%, 50%, 40%, and 100% for Diffusion Policy, Octo, OpenVLA, and $\pi_{0}$, respectively. But as shown in Table I, the performance drops to 0% of their real-world performance when evaluated in the real-to-sim setting for all models except for $\pi_{0}$. With ARRO, OpenVLA retains 55% of its real-world performance. Octo does not benefit as much and only achieves 5% of its original success rate. Diffusion Policy does not exhibit any immediate transfer in terms of success rate, which can be attributed, at least in part, to the absence of a pre-training stage and differences in the vision backbone architecture. $\pi_{0}$ already has a high baseline performance and can retain 94% of its success rate when masking is applied. The low relative performance of ARRO on $\pi_{0}$ can be explained by a high real-to-real success probability of 90%, while the real-to-sim success rate is comparable to that of the masked variant. In addition to the success rates in the real-to-sim experiments, we computed reward values using a simple distance metric inspired by ManiSkill [maniskill]. The results are shown in Fig. 9. For Diffusion Policy, Octo, and OpenVLA, both ARRO and the simple black background yield higher rewards than the vanilla baselines.
@@ IV-C1 Real-to-Sim
Figure 6: Performance comparison across three generalist models, Octo and OpenVLA and $\pi_{0}$, evaluated under altered visual conditions.
Figure 7: Wrist camera view of ARRO for the $\pi_{0}$ experiment “pick up the green cube”. Top: Example input image of a cluttered scene and the masked version. Bottom: Scene with a red cube that is masked by ARRO.
@@
Figure 8: Experimental setup for the pick task in two environments and one cross-embodiment scenario. Left to right: The real-world FR3 setup, the model of the setup in MuJoCo used for collecting sim-pick-v1, and the UR5e setup for cross-embodiment experiments in MuJoCo.
TABLE I: Relative policy performance of real-to-sim and cross-embodiment experiments after domain transfer.
@@
Policy Camera Control Dataset Vanilla Masked ARRO Dataset Vanilla Masked ARRO
Diffusion Policy Side Cartesian pick-v2 0% 0% 0% sim-pick-v1 0% 80% 99%
Octo Side Cartesian pick-v2 0% 3% 5% sim-pick-v1 71% 87% 87%
OpenVLA Side Cartesian pick-v2 0% 53% 55% sim-pick-v1 29% 72% 80%
$\pi_{0}$ Side + Wrist Joints pick-wrist-v1 79% 94% 77%
@@
Figure 9: Normalized mean per step reward over 100 episodes on the sim-pick-v1 task in simulation on the FR3 (same embodiment). All models were trained only on real-world datasets. Shaded areas indicate standard deviations.
IV-C2 Cross-Embodiment
To test the cross-embodiment capabilities of ARRO, we train our policies on a simulation dataset, referred to as sim-pick-v1, using the FR3 and test it on a UR5e embodiment. We evaluate both in-distribution on the training environment for 100 episodes and on the new UR5e embodiment for 100 episodes. The relative performance is shown in Table I under cross-embodiment. Note that $\pi_{0}$ is not included, as it operates in joint space and therefore cannot be directly applied to the UR5e robot due to its different kinematics.
@@ IV-C2 Cross-Embodiment
The diffusion policy’s performance drops to 0% when evaluated on a novel embodiment, which is expected given that it was trained exclusively on the FR3 robot. In contrast, both the masking with a black background and ARRO exhibit only a minor reduction in success rates, retaining most of their performance. This is likely because the embodiment change has minimal impact on the diffusion policy: the visual input remains largely unaffected due to the masking of the robot body, and the control commands are issued in absolute Cartesian space, which remains invariant to the embodiment.
@@
Octo and OpenVLA also exhibit reduced drops in success rates when combined with ARRO and masking. Unlike the Diffusion Policy, their performance does not drop to 0% in the vanilla setting for the new embodiment. This can be attributed to their larger and more diverse pre-training datasets, which are explicitly designed to support cross-embodiment generalization. As a result, both models encountered the UR5e robot during training, facilitating better transfer even in simulation settings.
V Limitations
While our method shows clear advantages across all evaluated models and tasks, its performance depends heavily on the underlying segmentation model, especially for wrist cameras, where objects can appear or disappear abruptly. In our experiments, task-relevant objects were occasionally not tracked reliably, leading to degraded performance. This issue could be mitigated by periodically re-initializing the segmentation model or by running a lightweight object detector in parallel to trigger tracker resets for the segmentation model as objects enter or leave the scene. Notably, our approach can still function under minor segmentation errors, depending on the visuomotor policy’s tolerance to such perturbations.
Finally, our method cannot compensate for reflections and changes in lighting. These types of domain shifts could be mitigated with image inpainting methods or integrating depth images. However, the real-to-sim results indicate that more recent models, such as $\pi_{0}$, are less sensitive to such changes.
VI Conclusion
@@
We introduce ARRO, a calibration-free visual augmentation pipeline that enhances the robustness of visuomotor policies by segmenting task-relevant regions and compositing them onto a structured virtual background. Our method improves generalization without requiring retraining or camera calibration, and experiments in both real and simulated settings show consistent gains for task-specific as well as generalist policies.
A possible direction for future work is to explore whether, for the wrist-mounted camera, replacing the static background overlay with dynamic, view-dependent frames would further enhance the results. This remains an open question and a promising area for future investigation.
@@ ACKNOWLEDGMENT
The authors acknowledge the HPC resources provided by the Erlangen National HPC Center (NHR@FAU) under the BayernKI project no. v106be.